OpenStudy (znappydooz):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 4, side BC measures x+8, and side CA measures 8. In triangle DEF, side DE measures 4, side EF measures 3x+3, and side FD measures 8. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (znappydooz):

@sleepyjess

OpenStudy (sleepyjess):

O_O

OpenStudy (znappydooz):

Its really annoying isnt it? XD

OpenStudy (znappydooz):

I hate it :/

OpenStudy (sleepyjess):

Ok, 

OpenStudy (sleepyjess):

This is the information we are given.

OpenStudy (sleepyjess):

We know that they are congruent, so x + 8 = 3x + 3

OpenStudy (znappydooz):

Oke..?

OpenStudy (sleepyjess):

Do you know how to solve for x?

OpenStudy (znappydooz):

I'm not so sure... can you guide me through? Do you have to try to leave x alone?

OpenStudy (sleepyjess):

We need to get all x's on one side, and all constants on the other side.

OpenStudy (sleepyjess):

So we start with x + 8 = 3x + 3. Subtract 3 from both sides

OpenStudy (znappydooz):

That would be x + 5 = x?

OpenStudy (sleepyjess):

Close, good try. We still have 3x on the right side though.

OpenStudy (sleepyjess):

So now we have x + 5 = 3x. Subtract x from both sides :)

OpenStudy (znappydooz):

5 = 3?

OpenStudy (znappydooz):

@cwrw238

OpenStudy (znappydooz):

@Straybullet, can u help by any chance?

OpenStudy (anonymous):

Okke so you have to remember that when youre subtracting youre only subtracting a constant, so when you subtract 3 from each side you are subtracting only the 3

OpenStudy (anonymous):

going from 3x + 3 to x would mean you hd divided by 3, not subtracted...so in order to subtract you merely take away the 3 and thats it

OpenStudy (znappydooz):

Sleepyjess explains some... but we stayed at 5 = 3

OpenStudy (anonymous):

okke so you have x + 5 = 3x The same thing applies...youre not actuatlly taking away the x entirely

OpenStudy (anonymous):

so on the one side, x - x = what?

OpenStudy (znappydooz):

Still a bit confused...

OpenStudy (anonymous):

Okke thats what im trying to explain so... thing of each part as totally separate

OpenStudy (anonymous):

on one side, you have (x) and (+5) and you are 'adding' (-x) the x's cancel each other out so you have just (5) on the one side

OpenStudy (znappydooz):

Would that then leave 5 = x?

OpenStudy (anonymous):

on the other side you have (3x) which is the same as saying that you have (x) (x) (x) or three xs and since you subtracted one x from the other side you only subject one x from this side

OpenStudy (anonymous):

Which means if you have: (x) (x) (x) - (x) what will you have?

OpenStudy (znappydooz):

So, what would I have left?

OpenStudy (znappydooz):

@sleepyjess, ur back :P

OpenStudy (anonymous):

Thats what im asking you, its very easy....you have three xs and you take away one

OpenStudy (sleepyjess):

If you have 3 kittens, and you give away 1 kitten, how many kittens do you have?

OpenStudy (znappydooz):

2...

OpenStudy (sleepyjess):

Yes, so we have 5 = 2x

OpenStudy (anonymous):

so 2x..... 5 = 2x

OpenStudy (znappydooz):

OOOHHH! I see what's going on now :P

OpenStudy (sleepyjess):

Now, divide each side by 2.

OpenStudy (znappydooz):

2.5 = x

OpenStudy (anonymous):

Yiiisss :E

OpenStudy (sleepyjess):

Yayz!

OpenStudy (znappydooz):

So, wait, what do I have to write down then?

OpenStudy (sleepyjess):

Just write down the process of going from x + 8 = 3x + 3 to x = 2.5

OpenStudy (anonymous):

So now you have to substitute 2.5 for x into one of the original equations

OpenStudy (sleepyjess):

Then do what @Straybullet said, substitute in 2.5

OpenStudy (anonymous):

Right how you solved it and then the final step is putting 2.5 as x in one of the equations, (x +8) would be the easier one

OpenStudy (znappydooz):

I'm confused again... sorry :/

OpenStudy (sleepyjess):

Do you know the steps to get from x + 8 = 3x + 3?

OpenStudy (znappydooz):

x + 8 = 3x + 3 x + 8 - 3 = 3x + 3 - 3 x + 5 = 3x

OpenStudy (sleepyjess):

Then what?

OpenStudy (znappydooz):

DIVIDE XD

OpenStudy (sleepyjess):

Subtract

OpenStudy (sleepyjess):

You forgot to give away the kitten

OpenStudy (anonymous):

Its subtract again cause you still have to subtract the one x from each side

OpenStudy (znappydooz):

OHH, okee

OpenStudy (sleepyjess):

After that we get 5 = 2x, then divide

OpenStudy (znappydooz):

Solving for x: x + 8 = 3x + 3 x + 8 - 3 = 3x + 3 - 3 x + 5 = 3x x - x + 5 = 3x - x 5 = 2x

OpenStudy (znappydooz):

divide by 2, and we get 2.5 = x?

OpenStudy (sleepyjess):

yep

OpenStudy (znappydooz):

Alright, now what comes next?

OpenStudy (sleepyjess):

Substitute in 2.5 for x

OpenStudy (znappydooz):

So, Line BC = 10.5, Line EF = 10.5

OpenStudy (znappydooz):

@sleepyjess, @Straybullet

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

Yiiss

OpenStudy (znappydooz):

THis is what I got written so far...

OpenStudy (znappydooz):

Solving for x: x + 8 = 3x + 3 x + 8 - 3 = 3x + 3 - 3 x + 5 = 3x x - x + 5 = 3x - x 5 = 2x 5/2 = 2x/2 2.5 = x Line BC = 10.5 Line EF = 10.5

OpenStudy (znappydooz):

Do I have to write something else?

OpenStudy (sleepyjess):

Show the work for getting to 10.g

OpenStudy (sleepyjess):

*10.5

OpenStudy (znappydooz):

What about these 2? What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

Is the equation you used in the beginning, the x + 8 = 3x + 3

OpenStudy (znappydooz):

Yes, i believe so :)

OpenStudy (anonymous):

I meant 'its' the equation that would help you to solve for the side length of BC and EF is x + 8 = 3x + 3